• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 02:42:57 on April13,2008

    Entry number 231652

    keyword=Beam alignment studies finished

    Beam alignment studies finished.

    We had real trouble getting the beam aligned. It is very unstable, and there are problems delivering high current to Hall C which seem to be affecting us.

    We checked position on BeO, asked MCC to restore nominal values.

    Requested a harp scan. IHA1H03A was not working well. On 1H03B we had sigmas around 100 microns in x and y.

    We then checked the holey carbon T1 and T2 rates against empty target rates at 5 microamps.

    Empty - T1 15000 Hz T2 19000 Hz
    Holey Carbon - T1 14000 Hz T2 18000 Hz (beam position was slightly better for C than empty).

    We are now taking Ta calibration data.